# Exploring Knowledge Tracing in Tutor-Student Dialogues

### v1

@@ 4.2 LLMKT
Another feature that makes dialogue KT different than the regular KT task is that most dialogue turns (at least 58%) are associated with more than one KC; see Table 1 for details. In contrast, most existing KT methods are developed for questions that cover a single KC; solutions that can deal with the occasional multi-KC question in existing KT datasets include i) converting a single question-response pair with multiple KCs to multiple pairs for different “questions”, each covering a single KC, and ii) treating a unique combination of multiple KCs as a new KC. However, due to the prevalence of dialogue turns tagged with multiple KCs, we found that these solutions are not applicable to the KT-in-dialogues setting. Therefore, we resort to an “AND”-type model, where the student is likely to respond to a tutor-posed task correctly only if they master all KCs involved in the turn. Formally, we model the probability of a correct student response at a turn as the product of the KC mastery across ones involved in this turn, i.e.,
$P(y_{j}=1)=\textstyle\prod_{k=1}^{K}P(z_{jk}=1).$ (1)
@@
We now detail our proposed LLM-based KT method, which we refer to as LLMKT. One can, in theory, estimate student knowledge states $z_{jk}$ by simply providing the text of the dialogue up to $t_{j}$ to a pre-trained LLM, and prompting it to predict whether the student masters the KCs $c_{jk}$ for each $k\in[1,K]$. However, since LLMs are not pre-trained on the specific KT task, we cannot expect them to do well in this “zero-shot” setting. Therefore, we adopt a fine-tuning approach using an open-source, decoder-only LLM, such as Llama 3 [34], for the KT task. We model the estimated probability of a student mastery $\hat{z}_{jk}$ as
$\hat{z}_{jk}=P_{\theta}(z_{jk}=1)=\frac{\exp(v^{\mathcal{T}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))}{\exp(v^{\mathcal{T}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))+\exp(v^{\mathcal{F}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))},$ (2)
@@
where $\theta$ represents the model parameters, and $v^{\mathcal{T}}_{\theta}(\cdot)$ and $v^{\mathcal{F}}_{\theta}(\cdot)$ return the output logits of the LLM at the “True” and “False” vocabulary tokens, respectively. In other words, with i) the entire conversation history and ii) a textual prompt instructing the LLM to estimate the student’s current KC mastery level as input, we use an LLM to parameterize this estimate through two token generation probabilities. In practice, to improve model efficiency, we pack all KCs in $C_{j}$ into a single prompt; we use customized attention masks and position embeddings to ensure that KCs do not attend to each other and their mastery levels are estimated independently. We do not explicitly provide KCs or correctness labels for previous turns due to memory constraints, but observe that LLMs can infer them from the context.
@@
We fine-tune the LLM by maximizing the likelihood of all observed student response correctness labels, or equivalently, minimizing the binary cross entropy (BCE) loss between the ground truth correctness labels $y_{j}$ and the predicted student correctness probability $\hat{y}_{j}$ as
$\hat{y}_{j}=\textstyle\prod_{k=1}^{K}\hat{z}_{jk},\quad\mathcal{L}(d;\theta)=-\textstyle\sum_{j=1}^{M}y_{j}\log\hat{y}_{j}+(1-y_{j})\log(1-\hat{y}_{j}).$ (3)
@@
While this method is relatively simple, it has many advantages. First, because we start with a powerful pre-trained LLM, LLMKT is able to leverage existing textual analysis capabilities that are not available in other KT architectures. Second, we do not add any new parameters to the pre-trained model, likely reducing the need for large amounts of training data. Third, since LLMKT operates on textual content rather than question/KC embeddings that need to be learned from scratch, it should be more generalizable to new KCs that are not seen during training.
@@ 4.3 DKT-Sem
In addition to LLMKT, we experiment with an alternative, simpler KT method for dialogues, which we can be viewed as a strong baseline. We slightly modify the deep KT (DKT) model [41] structure to use semantic embeddings of the textual content in dialogues, resulting in a method we term DKT-Sem. We simply use a learnable projection matrix to map text embeddings of every dialogue turn to the latent knowledge space in DKT, and similarly predict masteries with a learnable projection from the latent space to the text embeddings of KCs. We believe that this simple method can result in better generalization across KCs and dialogues compared to DKT, which learns question and KC representations from scratch. Formally, we model KC mastery probability as
$\mathbf{t}_{j}=\operatorname{S-BERT}(t_{j}),\quad\mathbf{s}_{j}=\operatorname{S-BERT}(s_{j}),\quad\mathbf{c}_{j}=\textstyle\frac{1}{K}\sum_{k=1}^{K}\operatorname{S-BERT}(c_{jk}),\quad\mathbf{y}_{j}=\operatorname{emb}({y_{j}})$ (4)
$\mathbf{x}_{j}=\tanh([\mathbf{t}_{j};\mathbf{s}_{j};\mathbf{c}_{j}]\mathbf{W})\oplus\mathbf{y}_{j},\quad\mathbf{h}_{j}=\operatorname{LSTM}(\mathbf{x}_{\leq j}),\quad\hat{\mathbf{z}}_{j+1}=\sigma(\mathbf{h}_{j}\mathbf{B}\mathbf{C}^{T}).$ (5)
The input at each time step is created by first computing the sentence BERT (S-BERT) [44] embeddings of the teacher turn, $\mathbf{t}_{j}$, student turn, $\mathbf{s}_{j}$, and element-wise average embeddings of the KCs, $\mathbf{c}_{j}$, respectively. Following DKT, response correctness is mapped to a learnable embedding, $\mathbf{y}_{j}$, which is then added element-wise to a linearly transformed version (through $\mathbf{W}$) of the concatenated tutor turn/student turn/KC embeddings, resulting in the input vector $\mathbf{x}_{j}$. Finally, the vector of estimated KC masteries at the next turn, $\hat{\mathbf{z}}_{j+1}$, are computed via a bilinear projection, $\mathbf{B}$, between the LSTM hidden state, $\mathbf{h}_{j}$, and the matrix of all KC S-BERT embeddings, $\mathbf{C}$, followed by the sigmoid function. We train DKT-Sem using the same objective as LLMKT (Eq. 3). Due to the smaller model size, DKT-Sem is easier to train than LLMKT, especially when access to computational resources to support LLM fine-tuning is limited.
@@ 5 Experiments
5.1 Datasets
We experiment with two datasets, CoMTA and MathDial, to study the dialogue KT task and validate the proposed KT methods. We acknowledge that there are many other datasets available, such as the ones in [13, 52, 53] and ones not focusing on math education. We select CoMTA since it is recent and highly reflective of AI-student dialogues that are becoming prevalent, and we select MathDial since it is relatively large-scale. However, our framework should be applicable to other math tutoring dialogue datasets as well. We show statistics for these datasets in Table 1.
Dataset Num. Dia. Num. Labels Correct Num. KCs Avg. KCs/Dia. Avg. KCs/Turn Turns >1 KC
@@ 5.1 Datasets
The MathDial dataset [33] contains 2,848 dialogues between students simulated by GPT-3.5 and crowd workers role-playing as human tutors. Each dialogue is in the context of an incorrect “student” solution to a math problem, and the tutor’s goal is to help the student arrive at the correct solution by helping them identify and correct their errors. We use the original train/test split, containing 2,235/588 dialogues, respectively, after removing a total of 25 dialogues, where GPT-4o’s annotation of student response correctness and turn KCs failed.
@@ 5.2 KT Methods
We evaluate multiple popular deep learning-based KT methods: DKT [41], DKVMN [58], AKT [16], SAINT [7], and simpleKT [29], all sourced from the pyKT code repository [30]. There are two key differences between these existing KT methods and our proposed KT methods for dialogues: i) they learn new embeddings for questions, KCs, and response correctness without leveraging any textual information, and ii) they are designed to analyze questions that cover a single KC. Since most of the dialogue turns contain at least two KCs in the CoMTA and MathDial datasets, we convert each turn into $K$ “pseudo-turns” that are each associated with one of the KCs, $c_{jk}$; all such pseudo-turns are associated with the same correctness label, $y_{j}$, of that turn. This modification follows common practice in the KT literature [16, 29]. We predict a single $\hat{\mathbf{z}}_{j}$ vector for each turn pair, using the output of the most recent time step that does not take $y_{j}$ as input, to avoid leaking the correctness label across pseudo-turns. We note that predicting a vector instead of a single probability requires altering the size of the output projection for all methods except DKT. We train all methods using the same objective as LLMKT and DKT-Sem (Eq. 3).
Bayesian KT (BKT) [9] is a classical KT method that uses a hidden Markov model to learn interpretable, per-KC learning acquisition parameters and use them to predict changing mastery levels for a student. Since BKT is designed to model data where each question is only assigned a single KC, we similarly expand turns into “pseudo-turns” as described above. Due to the rigidity of BKT, we change the setup slightly compared to other methods. At train time, we train on each pseudo-turn label directly rather than aggregating predictions at the true turn level. At test time, we collect predictions on all pseudo-turns and take their mean to get the predicted correctness for a real turn; taking a product, as the other methods do, would yield very low probabilities because BKT does not perform a product at train-time. In contrast to the other methods, this setup of predicting labels at the pseudo-turn level does not leak correctness labels because information does not interact across KCs in BKT. We use pyBKT [2] to train and generate predictions.
@@ 5.3 Metrics and Experimental Setup
To quantitatively evaluate the effectiveness of KT methods in the dialogue KT task, we employ 3 popular metrics: i) accuracy (Acc.), the portion of predicted labels that match the ground truth after rounding $\hat{y}_{j}$ to 0 or 1, ii) the area under the receiver operating characteristic curve (AUC), the most widely used metric in the KT literature, and iii) F1, the harmonic mean between binary precision and recall, again using rounded predictions. We compute these metrics on all correctness labels except for the first label in each dialogue, since existing KT methods are unable to make educated estimates without prior labels in context.
We implement LLMKT using the Llama-3.1-8B-Instruct model [34] in the Huggingface Transformers library [54] and adapt parameters in fine-tuning with LoRA [20]. We implement DKT-Sem using PyTorch and use the all-mpnet-base-v2 model to generate S-BERT embeddings. During training, we randomly take 20% of the dialogues in the train set to use as a validation set and perform early stopping by calculating the loss on this set. We then evaluate all metrics on the test set, and for CoMTA repeat training and testing for all 5 folds and report the average and standard deviation of each metric across folds. We train on a single NVIDIA RTX A6000 GPU.
Before computing final results, we perform a hyperparameter search to find optimal settings for each method. We perform an iterative search for LLMKT and a grid search for other methods, keeping the checkpoint with the highest AUC on the validation set. For LLMKT, we perform the search on the CoMTA dataset and use the same parameters on MathDial to reduce training time, and for other methods we use the best hyperparameter values found for each dataset. For LLMKT, we search over learning rate $\in\{5\cdot 10^{-5},1\cdot 10^{-4},2\cdot 10^{-4},3\cdot 10^{-4}\}$ and LoRA’s rank $r\in\{4,8,16,32\}$, set batch size to 64 using gradient accumulation, gradient norm clipping to 1.0, and train for 5 epochs. For other KT methods, we search over learning rate $\in\{1\cdot 10^{-4},2\cdot 10^{-4},5\cdot 10^{-4},1\cdot 10^{-3},2\cdot 10^{-3},5\cdot 10^{-3}\}$ and embedding size $\in\{8,16,32,64,128,256\}$, set batch size to 64, train for 100 epochs, and set all other model-specific hyperparameters to the defaults in pyKT. For all methods, we train with the AdamW optimizer with a weight decay of $1\cdot 10^{-2}$. We found the optimal learning rate and $r$ of LLMKT to be $2\cdot 10^{-4}$ and 16, respectively; the optimal hyperparameters for existing KT methods varied across methods and datasets, but tended to have smaller embedding sizes on CoMTA and larger ones on MathDial since the former is much smaller in scale compared to the latter.
@@ 6.1 Quantitative Results
Table 2 shows the quantitative performance of all KT methods on both datasets. We see that LLMKT significantly outperforms all other methods on both datasets, confirming its effectiveness in the dialogue setting. This advantage over other KT methods is likely due to the use of a pre-trained LLM to capture the rich textual content in dialogues. As a result, the LLM can estimate student knowledge after fine-tuning on student turn response correctness labels. We also see that DKT-Sem performs as well or better than existing deep learning KT methods even though it largely uses the same model architecture as DKT. This result further emphasizes the benefit of utilizing the textual dialogue context in some way, even if the original model architecture of existing KT methods remains largely unchanged.
Table 2: Student dialogue turn correctness prediction performance for all KT methods on both datasets, with best numbers in bold and second-best underlined. LLMKT significantly outperforms all other methods on both datasets.
@@
DKT $46.76_{\pm 5.59}$ $50.70_{\pm 6.33}$ $32.52_{\pm 9.48}$ 57.18 60.13 53.45
DKVMN $46.78_{\pm 4.76}$ $53.37_{\pm 7.84}$ $28.30_{\pm 3.70}$ 57.58 60.43 57.82
AKT $48.06_{\pm 3.48}$ $47.94_{\pm 6.50}$ $40.05_{\pm 8.73}$ 58.74 59.81 51.91
SAINT $48.71_{\pm 7.87}$ $47.00_{\pm 9.01}$ $32.30_{\pm 7.47}$ 54.96 58.01 51.57
simpleKT $48.53_{\pm 4.12}$ $47.64_{\pm 5.22}$ $42.54_{\pm 5.22}$ 57.38 60.50 53.97
DKT-Sem $\underline{58.59}_{\pm 8.64}$ $\underline{62.65}_{\pm 4.49}$ $\underline{60.93}_{\pm 14.22}$ 57.78 61.13 57.07
LLMKT $\textbf{60.36}_{\pm 5.95}$ $\textbf{65.15}_{\pm 6.62}$ $\textbf{65.54}_{\pm 6.96}$ 65.09 70.71 63.89
@@
On CoMTA, we see that all existing KT methods almost completely fail, unable to outperform the simple baseline of predicting the majority class label (which corresponds to an Acc. of $57.83_{\pm 5.08}$, AUC of $50.0_{\pm 0.00}$, and F1 of $58.87_{\pm 29.70}$). This result is expected since the size of the CoMTA dataset is small and neural network-based KT methods are data-hungry. Another important factor is that there are 164 unique KCs across only 623 labeled dialogue turn pairs, which is not enough for these methods to properly learn the KC embeddings or for BKT to learn the per-KC parameters. In contrast, LLMKT needs only minimal training data to reach acceptable accuracy, due to the underlying LLM being highly effective at understanding previous student dialogue turns. It can also leverage KC text information to generalize across KCs, which is especially beneficial for ones that do not occur frequently.
@@
On MathDial, which has 21x more data than CoMTA, we see that existing KT methods perform significantly better (for reference, the majority baseline only gets an Acc. of $52.64$ since class labels are balanced). This result is expected and shows that they can perform the dialogue KT task when given enough training data. BKT outperforms neural network-based KT methods since even the larger MathDial dataset has short sequence length (less than 5 turns on average). Therefore, recurrent or attention neural networks that can capture long context history may not outperform simple models in this setting. We also see that DKT-Sem does not outperform existing KT methods, which suggests that the way it uses textual content only helps when little training data is given. However, LLMKT still maintains a large advantage over all other KT methods. This result suggests that even on large-scale datasets, prior dialogue history can still significantly improve KT performance, even when labeled student turn correctness data is abundant. However, this rich contextual information needs to be leveraged effectively, likely by an LLM-based KT method.
@@
Finally, we see that performance of all KT methods is relatively low on the dialogue KT task, with a maximum of around $71\%$ AUC for student dialogue turn correctness prediction. This observation is in stark contrast to results on the standard KT task, where state-of-the-art KT methods can reach $80\%$ AUC for question response correctness prediction. This result suggests that dialogue KT is a challenging task, due to various reasons: First, the dialogues are relatively short, with only 4-5 labeled turn pairs on average, giving very little context for KT methods to leverage. For example, dialogues in CoMTA are cut short on purpose, to serve the goal in [35] on evaluating how accurate LLMs are at math. Therefore, we call upon researchers who have access to dialogues with more turns to validate our findings on their (possibly proprietary) data using our publicly available implementation. Second, student behavior in dialogues is more unpredictable compared to traditional KT, since there can be many diverse types of student responses to tutor turns. We believe that these challenges make dialogue KT a task worth studying in future work; it is likely that new KT methods need to be developed to improve performance on this task.
6.2 Qualitative Analysis

### v2

@@ 4.2 LLMKT
Another feature that makes dialogueKT different than the regular KT task is that most dialogue turns (at least 58%) are associated with more than one KC; see Table 1 for details. In contrast, most existing KT methods are developed for questions on a single KC; solutions that can deal with the occasional multi-KC question include i) converting a single question-response pair with multiple KCs to multiple pairs for different “questions”, each covering a single KC, and ii) treating a unique combination of multiple KCs as a new KC. However, due to the prevalence of dialogue turns tagged with multiple KCs, we found that these solutions are not applicable to our setting. Therefore, we consider the student’s likelihood of responding correctly to a tutor-posed task to be their average mastery of all KCs involved in the turn. Empirically, we found that averaging over KC masteries performed better than taking a product over them, consistent with findings in prior work [37]. Formally, we model the probability of a correct student response at a turn as
$P(y_{j}=1)=\textstyle\frac{1}{K}\sum_{k=1}^{K}P(z_{jk}=1).$ (1)
@@
In other words, we use a compensatory model rather than a conjunctive model [37]. We now detail our proposed LLM-based KT method, which we refer to as LLMKT. One can, in theory, estimate a student knowledge state, $z_{jk}$, by simply providing the text of the dialogue up to $t_{j}$ to a pre-trained LLM and prompting it to predict whether the student masters the KC, $c_{jk}$, for each $k\in[1,K]$. However, since LLMs are not pre-trained on the specific KT task, we cannot expect them to do well in this “zero-shot” setting. Therefore, we adopt a fine-tuning approach using Llama-3.1-8B-Instruct, an open-source, decoder-only LLM, for the KT task. We model the estimated probability of a student mastery $\hat{z}_{jk}$ as
$\hat{z}_{jk}=P_{\theta}(z_{jk}=1)=\frac{\exp(v^{\mathcal{T}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))}{\textstyle\sum_{\text{tok}\in\{\mathcal{T},\mathcal{F}\}}\exp(v^{\text{tok}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))},$ (2)
@@
where $\theta$ represents the model parameters, and $v^{\mathcal{T}}_{\theta}(\cdot)$ and $v^{\mathcal{F}}_{\theta}(\cdot)$ return the output logits of the LLM for the “True” and “False” vocabulary tokens, respectively. In other words, with i) the entire conversation history and ii) a textual prompt instructing the LLM to estimate the student’s current KC mastery level as input, we use an LLM to parameterize this estimate through two token generation probabilities. In practice, to improve model efficiency, we pack all KCs in $C_{j}$ into a single prompt; we use customized attention masks and position embeddings to ensure that KCs do not attend to each other and their mastery levels are estimated independently. We do not explicitly provide KCs or correctness labels for previous turns due to memory constraints, but observe that LLMs can infer them from the context.
@@
We fine-tune the LLM by maximizing the likelihood of all observed student response correctness labels, or equivalently, minimizing the binary cross entropy loss between the ground truth and predicted correctness values, $y_{j}$ and $\hat{y}_{j}$, as
$\hat{y}_{j}=\textstyle\frac{1}{K}\sum_{k=1}^{K}\hat{z}_{jk}$ (3)
$\mathcal{L}(d;\theta)=-\textstyle\sum_{j=1}^{M}y_{j}\log\hat{y}_{j}+(1-y_{j})\log(1-\hat{y}_{j}).$ (4)
@@
While this method is relatively simple, it has many advantages. First, because we start with a powerful pre-trained LLM, LLMKT is able to leverage existing textual analysis capabilities that are not available in other KT architectures. Second, we do not add any new parameters to the pre-trained model, likely reducing the need for large amounts of training data. Third, since LLMKT operates on textual content rather than question/KC embeddings that need to be learned from scratch, it should be more generalizable to new KCs that are not seen during training. We note that LLMKT is similar to CLST [24], which trains an LLM to predict student correctness using a softmax over tokens (Eq. 2). However, unlike LLMKT, their method does not handle multiple KCs per item, which is critical to the dialogueKT task.
@@ 4.3 DKT-Sem
In addition to LLMKT, we experiment with an alternative, simpler KT method for dialogues, which can be viewed as a strong baseline. We slightly modify the deep KT (DKT) model [46] to use semantic embeddings of the textual content in dialogues, resulting in a method we term DKT-Sem. We map the text embeddings of every dialogue turn to the latent knowledge space in DKT as input, and map the text embeddings of KCs to the latent knowledge space to predict KC masteries. We believe that this simple method can result in better generalization across KCs and dialogues compared to DKT, which learns question and KC representations from scratch. Formally, we model KC mastery probability as
$\mathbf{t}_{j}=\operatorname{S-BERT}(t_{j}),\quad\mathbf{s}_{j}=\operatorname{S-BERT}(s_{j}),\quad\mathbf{c}_{j}=\textstyle\frac{1}{K}\sum_{k=1}^{K}\operatorname{S-BERT}(c_{jk})$ (5)
$\mathbf{y}_{j}=\operatorname{emb}({y_{j}}),\quad\mathbf{x}_{j}=\tanh([\mathbf{t}_{j};\mathbf{s}_{j};\mathbf{c}_{j}]\mathbf{W})\oplus\mathbf{y}_{j}$ (6)
$\mathbf{h}_{j}=\operatorname{LSTM}(\mathbf{x}_{\leq j}),\quad\hat{\mathbf{z}}_{j+1}=\sigma(\mathbf{h}_{j}\mathbf{B}\mathbf{C}^{T}).$ (7)
The input at each time step is created by first computing the sentence BERT (S-BERT) [49] embeddings of the teacher turn, $\mathbf{t}_{j}$, student turn, $\mathbf{s}_{j}$, and element-wise average embeddings of the KCs, $\mathbf{c}_{j}$, respectively. Following DKT, response correctness is mapped to a learnable embedding, $\mathbf{y}_{j}$, which is then added element-wise to a transformed version (through $\mathbf{W}$ and $\tanh$) of the concatenated tutor turn/student turn/KC embeddings, resulting in the input vector $\mathbf{x}_{j}$. Finally, the vector of estimated KC masteries at the next turn, $\hat{\mathbf{z}}_{j+1}$, are computed via a bilinear projection, $\mathbf{B}$, between the LSTM hidden state, $\mathbf{h}_{j}$, and the matrix of all KC S-BERT embeddings, $\mathbf{C}$, followed by the sigmoid function. We train DKT-Sem using the same objective as LLMKT (Eq. 4). Due to the smaller model size, DKT-Sem is easier to train than LLMKT, especially when access to computational resources to support LLM fine-tuning is limited.
@@ 5 Experiments
Dataset Num. Dia. Num. Labels Correct Num. KCs Avg. KCs/Dia. Avg. KCs/Turn Turns w/ >1 KC
@@
5.1 Datasets
We experiment with two datasets, CoMTA and MathDial, to study the dialogueKT task and validate the proposed KT methods. We select CoMTA since it is recent and highly reflective of AI-student dialogues that are becoming prevalent, and we select MathDial since it is relatively large-scale. We acknowledge that there are many other datasets available, such as the ones in [14, 57, 59] and ones not focusing on math education; our overall framework should be applicable to these as well. We show statistics for CoMTA and MathDial in Table 1.
@@ 5.1 Datasets
The MathDial dataset [36] contains 2,848 dialogues between students simulated by GPT-3.5 and crowd workers role-playing as human tutors. Each dialogue is in the context of an incorrect “student” solution to a math problem, and the tutor’s goal is to help the student arrive at the correct solution by helping them identify and correct their errors. We use the original train/test split, containing 2,235/588 dialogues, respectively, after removing a total of 25 dialogues where GPT-4o’s annotation of student response correctness and turn KCs failed.
@@ 5.2 KT Methods
We evaluate multiple popular deep learning-based KT methods: DKT [46], DKVMN [64], AKT [17], SAINT [9], and simpleKT [32], all sourced from the pyKT code repository [33]. There are two key differences between these existing KT methods and our proposed KT methods for dialogues: i) they learn new embeddings for questions, KCs, and response correctness without leveraging any textual information, and ii) they are designed to analyze questions that cover a single KC. Since most of the dialogue turns contain at least two KCs in the CoMTA and MathDial datasets, we convert each turn into $K$ “pseudo-turns” that are each associated with one of the KCs, $c_{jk}$; all such pseudo-turns are associated with the same correctness label, $y_{j}$, of that turn. This modification follows common practice in the KT literature [17, 32]. We predict a single $\hat{\mathbf{z}}_{j}$ vector for each turn pair, using the output of the most recent time step that does not take $y_{j}$ as input, to avoid leaking the correctness label across pseudo-turns. We note that predicting a vector instead of a single probability requires altering the size of the output projection for all methods except DKT. We train all methods using the same objective as LLMKT and DKT-Sem (Eq. 4). We also compare against Bayesian KT (BKT) [11], a classical KT method that uses a hidden Markov model to learn interpretable, per-KC learning acquisition parameters and use them to predict changing mastery levels for a student. Since BKT is designed to model data where each question is only assigned a single KC, we similarly expand turns into “pseudo-turns” as described above. Due to the rigidity of BKT, we use a slightly different setup compared to other methods. At train-time, we train on all pseudo-turn labels, rather than aggregating at the true turn level. At test-time, we take the average pseudo-turn prediction to be the true turn prediction. In contrast to the other methods, this pseudo-turn level prediction setup does not leak correctness labels because information does not interact across KCs in BKT. We use pyBKT [2] to train and generate predictions.
@@ 5.3 Metrics and Experimental Setup
To quantitatively evaluate the effectiveness of KT methods in the dialogueKT task, we employ 3 widely used metrics: i) accuracy (Acc.), the portion of predicted labels that match the ground truth after rounding $\hat{y}_{j}$ to 0 or 1, ii) the area under the receiver operating characteristic curve (AUC), the most widely used metric in the KT literature as it accounts for class imbalance and all prediction thresholds, and iii) F1, the harmonic mean between binary precision and recall. We compute these metrics on all correctness labels except for the first label in each dialogue, since existing KT methods are unable to make educated estimates without prior labels in context.
We implement LLMKT using the Huggingface Transformers library [60] and perform fine-tuning with LoRA [22] on NVIDIA RTX A6000 GPUs. We implement DKT-Sem using PyTorch and use the all-mpnet-base-v2 S-BERT model to generate text embeddings. During training, we randomly take 20% of the dialogues in the train set to use as a validation set and perform early stopping. We then evaluate all metrics on the test set, and repeat for all 5 folds on CoMTA.
Before computing final results, we perform a grid search to find optimal hyperparameters for each method, keeping the checkpoint with the highest AUC on the validation set. For LLMKT, we perform the search on the CoMTA dataset and use the same parameters on MathDial to reduce training time, and for other methods we use the best hyperparameter values found for each dataset. For LLMKT, we search over learning rate $\in\{5\cdot 10^{-5},1\cdot 10^{-4},2\cdot 10^{-4},3\cdot 10^{-4}\}$ and LoRA’s rank $r\in\{4,8,16,32\}$, and fix LoRA’s $\alpha$ to 16, batch size to 64 using gradient accumulation, gradient norm clipping to 1.0, and train for 5 epochs. For other KT methods, we search over learning rate $\in\{1\cdot 10^{-4},2\cdot 10^{-4},5\cdot 10^{-4},1\cdot 10^{-3},2\cdot 10^{-3},5\cdot 10^{-3}\}$ and embedding size $\in\{8,16,32,64,128,256,512\}$, and fix batch size to 64, train for 100 epochs, and set all other model-specific hyperparameters to the defaults in pyKT. For all methods, we train with the AdamW optimizer with a weight decay of $1\cdot 10^{-2}$. We found the optimal learning rate and $r$ of LLMKT to be $2\cdot 10^{-4}$ and 16, respectively; the optimal hyperparameters for other KT methods varied across methods and datasets, but tended to have smaller embedding sizes on CoMTA and larger ones on MathDial since the former is much smaller in scale compared to the latter.
@@ 6.1 Quantitative Results
DKT $52.59_{\pm 7.37}$ $53.20_{\pm 10.50}$ $49.56_{\pm 4.73}$ $59.90$ $63.22$ $55.18$
DKVMN $47.22_{\pm 9.12}$ $46.81_{\pm 8.04}$ $48.80_{\pm 12.37}$ $57.53$ $60.42$ $52.45$
AKT $49.95_{\pm 4.74}$ $51.40_{\pm 7.34}$ $50.78_{\pm 11.04}$ $62.07$ $63.31$ $57.14$
SAINT $49.17_{\pm 4.88}$ $47.81_{\pm 3.89}$ $50.38_{\pm 6.68}$ $55.97$ $60.10$ $54.24$
simpleKT $51.25_{\pm 4.05}$ $51.25_{\pm 5.60}$ $49.82_{\pm 10.45}$ $59.24$ $63.83$ $53.85$
DKT-Sem $\underline{56.83}_{\pm 9.75}$ $\underline{61.82}_{\pm 6.56}$ $\mathbf{63.27}_{\pm 9.57}$ $\underline{62.17}$ $\underline{66.18}$ $\underline{59.30}$
LLMKT $\mathbf{58.01}_{\pm 5.40}$ $\mathbf{65.79}_{\pm 4.44}$ $\underline{60.72}_{\pm 10.77}$ $\mathbf{68.41}$ $\mathbf{76.71}$ $\mathbf{62.21}$
Table 2: Student dialogue turn correctness prediction performance (averaged with standard deviation shown when applicable) for all KT methods, with best numbers in bold and second-best underlined. LLMKT outperforms all other methods on both datasets.
Table 2 shows the quantitative performance of all KT methods on both datasets. We see that LLMKT significantly outperforms all existing KT methods and generally outperforms DKT-Sem on both datasets, confirming its effectiveness in the dialogue setting. This advantage over other KT methods is likely due to the use of a pre-trained LLM to capture the rich textual content in dialogues. As a result, the LLM can estimate student knowledge after fine-tuning on student turn response correctness labels. We also see that DKT-Sem performs better than all existing KT methods, even though it largely uses the same model architecture as DKT. This result further emphasizes the benefit of utilizing the textual dialogue context in some way, even if the original model architecture of existing KT methods remains largely unchanged.
@@
On CoMTA, we see that all existing KT methods almost completely fail, unable to outperform the simple baseline of predicting the majority class label (which corresponds to an Acc. of $57.83_{\pm 5.08}$, AUC of $50.0_{\pm 0.00}$, and F1 of $58.87_{\pm 29.70}$). This result is expected since the size of the CoMTA dataset is small and neural network-based KT methods are data-hungry. Another important factor is that there are 164 unique KCs across only 623 labeled dialogue turn pairs, which is not enough for these methods to properly learn the KC embeddings or for BKT to learn the per-KC parameters. In contrast, LLMKT needs only minimal training data to reach acceptable accuracy, due to the underlying LLM being highly effective at understanding previous student dialogue turns. It can also leverage the textual content of the KC definitions to generalize across KCs, which is especially beneficial for ones that do not occur frequently.
@@
On MathDial, which has 21x more data than CoMTA, we see that existing KT methods perform significantly better (for reference, the majority baseline only gets an Acc. of $52.64$ since class labels are balanced). This result is expected and shows that they can perform the dialogueKT task when given enough training data. While DKT-Sem still outperforms existing KT methods, it does so by a smaller margin than on CoMTA, indicating DKT-Sem has its greatest advantage when little training data is given. However, LLMKT still maintains a large advantage over all other KT methods. This result suggests that even on large-scale datasets, LLM inference over prior dialogue history and textual content of KCs can still significantly improve KT performance, even when labeled student turn correctness data is abundant.
@@
Finally, we see that performance of all KT methods is relatively low on the dialogueKT task, with a maximum of around $76\%$ AUC for student dialogue turn correctness prediction. This result is in stark contrast to results on the standard KT task, where state-of-the-art KT methods can surpass $80\%$ AUC for question response correctness prediction on some datasets. This observation suggests that dialogueKT is a challenging task, due to various reasons: First, dialogues can be relatively short, giving very little context for KT methods to leverage. For example, dialogues in CoMTA are cut short on purpose, with only 4-5 labeled turn pairs on average, to serve the goal in [39] on evaluating how accurate LLMs are at math. Therefore, we call upon researchers who have access to dialogues with more turns to study these methods on their (possibly proprietary) data using our publicly available implementation. Second, student behavior in dialogues is more unpredictable compared to traditional KT, since there can be many diverse types of student responses to tutor turns. We believe that these challenges make dialogueKT a task worth studying in future work; it is likely that new KT methods need to be developed to improve performance on this task.
6.2 Qualitative Analysis: Case Study
Speaker Turn Text KCs and Predicted Masteries
Tutor Point X is at (-4, 6) and point Z is at (6, -2). Determine the coordinates of point Y on line segment XZ such that XY is 3 times as long as YZ. What do you think is the important information here? 1) Use a pair of perpendicular number lines to define a coordinate system. $0.4688$
2) Apply the Pythagorean Theorem to find the distance between two points. $0.4688$
3) Recognize and represent proportional relationships between quantities. $0.5622$
4) Find the point on a line segment between two points that partitions the segment in a given ratio. $0.6225$
Student Firstly, I need to determine the distance between point X and Z. That would be the square root of 164, which is roughly 12.8
Tutor Excellent beginning! You’ve calculated the total length of line segment XZ. Now, how can we use this information to determine the coordinates of point Y? Keep in mind, XY is 3 times as long as YZ. 1) Recognize and represent proportional relationships between quantities. $0.3208$
2) Find the point on a line segment between two points that partitions the segment in a given ratio. $0.3486$
Student I would divide 12.8 by 3 total parts.
Table 3: An example dialogue with student turn correctness labels, KC labels, and LLMKT’s predicted mastery level for each KC. LLMKT accurately predicts student correctness at both turns by leveraging the difficulty of the tutor-posed questions.